Construct the confidence interval for the ratio of the population variances given the following sample statistics. Round your answers to four decimal places. n1=12 , n2=19, s12=8.041, s22=4.964, 90% level of confidence
Answer #1

At 90% confidence level, the critical values are

FL = F(0.95, 11, 18) = 0.3744

FR = F(0.05, 11, 18) = 2.3742

The 90% confidence interval for \frac{\sigma_{1}^2}{\sigma_{2}^2} is

\frac{s_{1}^2}{s{_2}^2} * \frac{1}{F_{R}} < \frac{\sigma_{1}^2}{\sigma_{2}^2} < \frac{s_{1}^2}{s{_2}^2} * \frac{1}{F_{L}}

= \frac{8.041}{4.964} * \frac{1}{2.3742} < \frac{\sigma_{1}^2}{\sigma_{2}^2} < \frac{8.041}{4.964} * \frac{1}{0.3744}

= 0.6823 < \frac{\sigma_{1}^2}{\sigma_{2}^2} < 4.3266